A rectangular painting covers an area of 12 square feet on a wall. The width of the painting is 1 foot more than its length.
Find the length and width of the painting.

Answer by fractalier(6550)

Area A = LW = 12.
Also we have W = L + 1.
Substituting we have
L(L + 1) = 12
L^2 + L - 12 = 0
Factoring we get
(L + 4)(L - 3) = 0
and the length is
L = 3
W = 4
(We could have seen this from the beginning but this is the algebra for it...)